# RelMap: Enhancing Online Map Construction with Class-Aware Spatial Relation and Semantic Priors

### v1

@@ Abstract
Online high-definition (HD) map construction plays an increasingly important role in scaling autonomous driving systems. Transformer-based methods have become prevalent in online HD map construction; however, existing approaches often neglect the inherent spatial and semantic relationships among map elements, which limits their accuracy and generalization. To address this, we propose RelMap, an end-to-end framework that enhances online map construction by incorporating spatial relations and semantic priors. We introduce a Class-aware Spatial Relation Prior, which explicitly encodes relative positional dependencies between map elements using a learnable class-aware relation encoder. Additionally, we propose a Mixture-of-Experts (MoE)-based Semantic Prior, which routes features to class-specific experts based on predicted class probabilities, refining instance feature decoding. Our method is compatible with both single-frame and temporal perception backbones, achieving state-of-the-art performance on both the nuScenes and Argoverse 2 datasets.
@@ 1 Introduction
Figure 1: Two major improvements proposed in RelMap for online vectorized map construction. The Class-aware Spatial Relation Prior explicitly encodes spatial dependencies between map elements, while the MoE-based Semantic Prior dynamically routes features to class-specific experts based on predicted map class probabilities for refined decoding.
@@ 3.2 Class-aware Spatial Relation Prior
Spatial Relation Encoder. Given a set of predicted map instances $\{(x^{l-1}_{i},y^{l-1}_{i})\}_{i=1}^{N_{p}}$ from the Transformer decoder layer $l-1$, we compute their relative spatial relationships using bounding box-based positions. To obtain the bounding box for each instance, we use the smallest axis-aligned bounding box that encloses all the $N_{p}$ points in the instance. Each instance is then represented by its bounding box $\mathbf{b}_{i}=(cx_{i},cy_{i},w_{i},h_{i})$, where $(cx_{i},cy_{i})$ denotes the center and $(w_{i},h_{i})$ represents the width and height. This provides a compact representation of the instance’s spatial extent without requiring additional model layers. To encode spatial dependencies, we construct a position relation embedding following [16] by computing the relative displacement and scale differences between instances. For each pair of map instances $i$ and $j$, we compute their relations:
@@ 4.3 Main Results
Quantitative Results on nuScenes. We evaluate both single-frame and temporal variants of our RelMap framework on the nuScenes [1] validation set. As shown in Tab. 1, RelMap-SF consistently achieves the highest mAP among single-frame methods. At 24 epochs, it reaches an mAP of 68.4, outperforming MapTRv2 by 6.9 points. At 110 epochs, it achieves an mAP of 74.0, exceeding MapQR [34] by 1.4 points and outperforming MapTRv2 by 5.3 points. These substantial gains underscore the efficacy of incorporating spatial and semantic priors. Compared to PriorMapNet [49], which relies on offline priors that require dataset-specific recomputation, our learnable priors within the Transformer model adapt dynamically during training, eliminating the need for precomputed priors while still achieving a 1.3-point higher mAP than PriorMapNet. For temporal map construction, RelMap-TF achieves a new state-of-the-art performance with an mAP of 77.1, outperforming leading methods including MapTracker [4], HisTrackMap [57], and MapExpert [60]. To better assess generalization capabilities, we further evaluate RelMap-TF on the geographically disjoint train/validation split proposed by StreamMapNet [58], which mitigates the risk of inflated scores due to memorization in the original split. As shown in Tab. 2, RelMap-TF achieves an mAP of 41.0, outperforming MapTracker by 0.7 points. This result demonstrates the model’s robustness and improved generalization to unseen environments.
@@ A.1 Class-aware Spatial Relation Prior
$d_{t}(i)=T^{\frac{2i}{D_{\text{PE}}}},\quad i=0,1,\dots,D_{\text{PE}}/2-1,$ (S2)
@@ C.1 Ablation Studies
(2) Four-point Representation: A map instance is represented by a set of four key points selected at fixed intervals along its sequence of points. Specifically, the four points $(x_{0},y_{0})$, $(x_{6},y_{6})$, $(x_{12},y_{12})$, and $(x_{19},y_{19})$ are chosen from the 202020 points that define each instance.

### v2

@@ Abstract
Online high-definition (HD) map construction is crucial for scaling autonomous driving systems. While Transformer-based methods have become prevalent in online HD map construction, most existing approaches overlook the inherent spatial dependencies and semantic relationships among map elements, which constrains their accuracy and generalization capabilities. To address this, we propose RelMap, an end-to-end framework that explicitly models both spatial relations and semantic priors to enhance online HD map construction. Specifically, we introduce a Class-aware Spatial Relation Prior, which explicitly encodes relative positional dependencies between map elements using a learnable class-aware relation encoder. Additionally, we design a Mixture-of-Experts-based Semantic Prior, which routes features to class-specific experts based on predicted class probabilities, refining instance feature decoding. RelMap is compatible with both single-frame and temporal perception backbones, achieving state-of-the-art performance on both the nuScenes and Argoverse 2 datasets.
@@ 1 Introduction
Figure 1: Two major improvements proposed in RelMap for online vectorized map construction. The Class-aware Spatial Relation Prior explicitly encodes spatial dependencies between map elements and decodes important weights, while the MoE-based Semantic Prior routes features to class-specific experts based on predicted class probabilities for refined decoding.
@@ 3.2 Class-aware Spatial Relation Prior
Spatial Relation Encoder. Given a set of predicted map instances $\{(x^{l-1}_{i},y^{l-1}_{i})\}_{i=1}^{N_{p}}$ from the Transformer decoder layer l1l-1, we compute their relative spatial relationships using bounding box-based positions. To obtain the bounding box for each instance, we use the smallest axis-aligned bounding box that encloses all the $N_{p}$ points in the instance. Each instance is then represented by its bounding box $\mathbf{b}_{i}=(cx_{i},cy_{i},w_{i},h_{i})$, where $(cx_{i},cy_{i})$ denotes the center and $(w_{i},h_{i})$ represents the width and height. This provides a compact representation of the instance’s spatial extent without requiring additional model layers. To encode spatial dependencies, we construct a position relation embedding following [16] by computing the relative displacement and scale differences between instances. For each pair of map instances $i$ and $j$, we compute their relations:
@@ 4.3 Main Results
Quantitative Results on nuScenes. We evaluate both single-frame and temporal variants of our RelMap framework on the nuScenes [1] validation set. As shown in Tab. 1, RelMap-SF consistently achieves the highest mAP among single-frame methods. At 24 epochs, it reaches an mAP of 68.4, outperforming MapTRv2 by 6.9 points. At 110 epochs, it achieves an mAP of 74.0, exceeding MapQR [34] by 1.4 points and outperforming MapTRv2 by 5.3 points. These substantial gains underscore the efficacy of incorporating spatial and semantic priors. Compared to PriorMapNet [49], which relies on offline priors that require dataset-specific recomputation, our learnable priors within the Transformer model adapt dynamically during training, eliminating the need for precomputed priors while still achieving a 1.3-point higher mAP than PriorMapNet. For temporal map construction, RelMap-TF achieves a new state-of-the-art performance with an mAP of 77.1, outperforming leading methods including MapTracker [4], HisTrackMap [57], and MapExpert [60]. To better assess generalization capabilities, we further evaluate RelMap-TF on the geographically disjoint train/validation split proposed by StreamMapNet [58], which mitigates the risk of inflated scores due to memorization in the original split. As shown in Tab. 2, RelMap-TF achieves an mAP of 42.3, demonstrating the model’s robustness and improved generalization to unseen environments.
@@ A.1 Class-aware Spatial Relation Prior
$d_{t}(i)=T^{\frac{2i}{D_{\text{PE}}}},\qquad i=0,1,\dots,D_{\text{PE}}/2-1,$ (S2)
@@ C.1 Ablation Studies
(2) Four-point Representation: A map instance is represented by a set of four key points selected at fixed intervals along its sequence of points. Specifically, the four points $(x_{0},y_{0})$, $(x_{6},y_{6})$, $(x_{12},y_{12})$, and $(x_{19},y_{19})$ are chosen from the $20$ points that define each instance.